[Brahms-l] Shift report 20050213 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 13 2005 - 16:05:15 EST
1. Came without beam, the machine has problem with power supply obtained beam at 11:30, but they dumped the beam at 15:50 before scheduled time(16:35)
2. Runs:
FS@4B2442, MRS@40A2255
RunNo - Trig2 Trig3
13161 - 146k , 70k
13162 - 266k , 110k

FS@4B2442, MRS@40B2255
13163 - 0k , 21k (MRS only)
13164 - 143k, 118k
13165 - 117k, 159k

Pedestals: 13159, 13160.
13168 - running.


3. No detector problems observed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 13 16:05:48 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 13 2005 - 16:05:57 EST